# Parse Trees Guided LLM Prompt Compression

### v1

@@ IV-F Algorithm and Complexity Analysis of PartPrompt
Input: Original Prompt $R$, Compression Ratio List $[c_{1},c_{2},\cdots,c_{u}]$, Node Value Adjustment Parameters $a_{1}$, $a_{2}$
Output: Compressed Prompts List of each Compression Ratio $[R_{cp,1},R_{cp,2},\cdots,R_{cp,u}]$
@@
3Sentence List $[R_{1},R_{2},\cdots,R_{m}]$, Structure Label List $[g_{1},g_{2},\cdots,g_{m}]$leftarrow PromptStructureParser($R$)
@@
5       Token List $[s_{1},s_{2},\cdots,s_{l_{j}}]_{j}$, Entropy List $[e_{1},e_{2},\cdots,e_{l_{j}}]_{j}$leftarrow SmallModel($R_{j}$)
6      Node List $[r_{1},r_{2},\cdots,r_{n_{j}}]_{j}$, Edge List $[(r_{w_{1}},r_{w_{2}}),(r_{w_{2}},r_{w_{3}}),\cdots,(r_{w_{h_{j}}},r_{w_{n_{j}}})]_{j}$leftarrow SentenceParser($R_{j}$)
7      Aligned Entropy List $[\tilde{e}_{1},\tilde{e}_{2},\cdots,\tilde{e}_{n_{j}}]_{j}$, Length List $[C_{1},C_{2},\cdots,C_{n_{j}}]_{j}\leftarrow$ TokenAlignment(Token List, Node List, Entropy List)
@@
10Adjusted Node Value Lists $[\hat{e}_{1},\hat{e}_{2},\cdots,\hat{e}_{n}]\leftarrow$ NodeValueAdjustment($T$, Aligned Entropy Lists, $a_{1}$, $a_{2}$)
11Solution $Q$leftarrow RecursiveTreeCompression($T$, Length Lists, Adjusted Node Value Lists)
12Compressed Prompts List $[R_{cp,1},R_{cp,2},\cdots,R_{cp,u}]$leftarrow TokenConcatenation(Solution, Compression Ratio List)
@@
[…]
@@ IV-G Theoretical Analysis
LLMLingua, another prominent selective prompt compression method, consists of the budget controller, iterative token-level prompt compression, and distribution alignment [13]. The budget controller divides the prompt into paragraphs, sorting them by information entropy and preserving those with higher entropy. In PartPrompt, the node value adjustment algorithm for actual nodes is defined as $E_{\text{Adjusted}}(v_{i})\leftarrow E_{\text{Aligned}}(v_{i})+M^{a_{1}}$ (Line 18 of Algorithm 1). Here $M$ is derived from the root-ward propagation and convergence of the information entropy of the corresponding segemts. In light of the fact that $M>1$ is typically the case, setting $a_{1}\gg 1$ gives rise to the conclusion that $M^{a_{1}}\gg E_{\text{Aligned}}(v_{i})$, which leads to the approximation that $E_{\text{Adjusted}}(v_{i})\approx M^{a_{1}}$. In this case, parse tree compression is equivalent to retaining the part with the larger value of $M$. Therefore, the budget controller of LLMLingua can be considered as a variant of parse tree compression which involves flattening the virtual part of the global parse tree and setting parameter $a_{1}\gg 1$ in the node value adjustment module of PartPrompt.
Considering $r_{j,i}$ as the $i$-th token of the $j$-th sentence, the iterative token-level prompt compression of LLMLingua employs Equation (11) to calculate the information entropy. This approach may be regarded as an intermediate stage between the initial version employed by Selective-Context, i.e., Equation (12), and the local approximation version of PartPrompt, i.e., Equation (13).
@@ V-A Datasets
Four representative datasets are employed to evaluate the proposed method. Concretely, BBCnews dataset is collected from articles on the BBC News website [32], and arXiv dataset is collected from papers on the arXiv preprint platform [33]. Following Selective-Context [12] and LLMLingua [13], these two datasets are used for testing contextual understanding prompts with the task of summarizing articles. To ensure that the model has not seen these data during training, we re-crawl new data for BBC News and arXiv, with all the release dates of these data occurring after Jan 1, 2024.
Due to the length of arXiv dataset, which exceeds the input length that Selective-Context and LLMLingua can handle [12, 13], we truncate the first 3000 tokens of the main text, referring as truncate arXiv. The original arXiv dataset is also used to verify the applicability of our method for texts exceeding the input limit of compared methods.
Regarding another two datasets, HotpotQA [34] is a multi-hop question answering dataset, which is designed to test the performance for question answering and multi-hop reasoning prompts. We use the same setting as in [35] and select 500 questions for testing. GSM8K [36] is a classic mathematical reasoning dataset. Following [13], we adopt the chain-of-thought prompt provided by [7] to evaluate the performance for in-context learning and chain-of-thought prompts.
@@ V-B Compared Methods for Prompt Compression
As our method belongs to selective prompt compression, we compare it to the state-of-the-art selective prompt compression methods. Selective-context [12] is the first to propose selective prompt compression method. It uses a smaller LLM to calculate the information entropy of tokens and removes words with lower information entropy. The question text is fully preserved for better performance. LLMLingua [13] further proposes the model distribution alignment, budget controller and token-level iterative compression algorithm based on Selecting-context. LLMLingua receives significant improvements in context-understanding prompts, and makes selective compression method viable for chain-of-thought prompts.
@@ V-D Evaluation
Following [12, 13], we take BLEU [39], Rouge (including Rouge1, Rouge2, RougeL) [40], and BERTScore (specifically BS-F1) [41] as the metric for evaluation on BBCnews and (truncate) arXiv datasets. BLEU is a composite metric of four metrics: 1-gram, 2-gram, 3-gram, and 4-gram, and is often used in the natural language processing. Rouge is classic linguistic metrics based on rule matching, while BERTScore calculates the similarity of the embeddings from BERT. Regarding above two datasets, the ground-truth answer is set to the output of the uncompressed prompt. For hotpotQA dataset, we follow [35] to compute precision, recall, and F1, since the answers is quite brief. For GSM8K dataset, as the questions are math-related, we adopt exact matching to calculate the EM score obeying the same setting as in [13, 7].
@@ V-E Other Settings
Apart from above settings, other experimental settings are clarified as follows. The information entropy is calculated using Llama2-7B [42], and the local parse tree is analyzed using the Stanford CoreNLP [31]. When adjusting the value in parse trees, the related hyper-parameters are set as follows: for BBCnews, $a_{1}=4$, $a_{2}=100$; for truncate arXiv, $a_{1}=3$, $a_{2}=100$; for hotpotQA, $a_{1}=3$, $a_{2}=20$; for GSM8K, $a_{1}=5$, $a_{2}=25$. And the typical range of them are $0\leq a_{1}\leq 5$, $1\leq a_{2}\leq 1000$. The maximum token length for target LLMs is set to 300 during inference for all experiments. For the pretrained language models used in this work, Mixtral-8x7B and Llama3-70B are accessed through HuggingFace Pro API [^3] Mixtral-8x7B and Llama3-70B: https://huggingface.co/blog/inference-pro; Qwen2-72B is accessed via Aliyun or can be downloaded from HuggingFace [^4] Qwen2-72B: https://huggingface.co/Qwen/Qwen2-72B-Instruct; Llama2-7B is downloaded from HuggingFace [^5] Llama2-7B: https://huggingface.co/meta-llama/Llama-2-7b and run on an NVIDIA GeForce RTX 3090 GPU with 24G memory.
@@
PartPrompt 13.69 43.54 22.87 41.05 67.58 173.7 4.96 11.16 39.17 17.96 36.36 66.99 556.6 5.38
30% token constraint
@@
methods hotpotQA GSM8K
BLEU Rouge1 Rouge2 RougeL BS-F1 precsion recall F1 Tokens $1/\tau$    EM Tokens $1/\tau$
20% token constraint
30% token constraint
50% token constraint
@@ VI Experiments
Extensive experiments are conducted to demonstrate the effectiveness and superiority of PartPrompt (the proposed method) regarding various prompt compression methods, datasets, metrics, compression ratios, and scenarios. Section VI-A presents a thorough comparison between PartPrompt and other compression methods to confirm the superiority of the proposed method. Considering the effect of compression ratios and target LLMs for inference, PartPrompt is further compared with other methods for various compression ratios in Section VI-B and different target LLMs in Section VI-C. Section VI-D provides a comprehensive ablation study to verify the positive role of each component in PartPrompt. Section VI-E investigates the potential of PartPrompt under the extreme long prompt scenario. In Section VI-F and VI-G, the compressed prompt is directly compared with uncompressed one; accordingly more abundant metrics and an intuitive example are employed to show the advantages of PartPrompt in terms of text similarity and coherence maintenance.
@@ VI-A Comparative Study: Main Experiments
Thorough experiments are conducted in this section to fully compare the proposed method with other selective compression methods. Four diverse datasets and multiple metrics are employed to evaluate the performance. For clarity, we choose three compression ratios for analysis: 20%, 30%, and 50%. As the actual compression ratio is not under the control of some methods, the average length of the compressed prompt (named as “Tokens”) and inverse compression ratio ($1/\tau$) are also listed to show the actual compression result. The LLM Generation is presented with a single compression ratio, since there is no compression ratio to set.
The experimental results are shown in Table I. Overall, it is evident that PartPrompt achieves remarkable improvements in almost all datasets, compression ratios, and metrics, which demonstrates the advantages of PartPrompt. In particular, the proposed PartPrompt receives the best performance compared to Selective-Context and LLMLingua, i.e., the two state-of-the-art selective prompt compression methods, on BBCnews, truncated arXiv, and hotpotQA datasets across all metrics and compression ratios. For the GSM8K dataset, PartPrompt is more robust than LLMLingua under different compression ratios and achieves the best averaged performance regarding the three compression ratios.
Apart from comparing to selective compression methods, PartPrompt is further compared to the generative compression method. Qwen2-72B is employed for generating compressed prompts, which is denoted as LLM Generation in Table I. Note that, using an LLM for compression needs a significant amount of computational resources, which is contrary to the original goal of prompt compression and therefore impractical. Despite that, LLM Generation may still serve as a reference method. The results show that PartPrompt outperforms LLM Generation on the BBCnews, truncated arXiv, and GSM8K datasets. This not only proves the advantage of PartPrompt but also validates the rationality of selective compression methods. For hotpotQA, LLM Generation obtains better performance than all selective compression methods, which may indicate the limitation of current selective compression methods in handling this kind of datasets with incohesive texts.

### v3

@@ IV-F Algorithm and Complexity Analysis of PartPrompt
Input: Original Prompt $R$, Compression Ratio List $\tau_{1},\tau_{2},\cdots,\tau_{u}]$, Node Value Adjustment Parameters $a_{1}$, $a_{2}$
Output: Compressed Prompts List of each Compression Ratio $R_{cp,1},R_{cp,2},\cdots,R_{cp,u}]$
@@
3Sentence List $R_{1},R_{2},\cdots,R_{m}]$, Structure Label List $g_{1},g_{2},\cdots,g_{m}] \leftarrow$ PromptStructureParser($R$)
@@
5   Token List $s_{1},s_{2},\cdots,s_{l_{j}}]_{j}$, Entropy List $e_{1},e_{2},\cdots,e_{l_{j}}]_{j} \leftarrow$ SmallModel($R_{j}$)
6  Node List $r_{1},r_{2},\cdots,r_{n_{j}}]_{j}$, Edge List $(r_{w_{1}},r_{w_{2}}),(r_{w_{2}},r_{w_{3}}),\cdots,(r_{w_{h_{j}}},r_{w_{n_{j}}})]_{j} \leftarrow$ SentenceParser($R_{j}$)
7  Aligned Entropy List $\tilde{e}_{1},\tilde{e}_{2},\cdots,\tilde{e}_{n_{j}}]_{j}$, Length List $C_{1},C_{2},\cdots,C_{n_{j}}]_{j}\leftarrow$ TokenAlignment(Token List, Node List, Entropy List)
@@
10Adjusted Node Value Lists $\hat{e}_{1},\hat{e}_{2},\cdots,\hat{e}_{n}]\leftarrow$ NodeValueAdjustment($T$, Aligned Entropy Lists, $a_{1}$, $a_{2}$)
11Solution $Q$ $\leftarrow$ RecursiveTreeCompression($T$, Length Lists, Adjusted Node Value Lists)
12Compressed Prompts List $R_{cp,1},R_{cp,2},\cdots,R_{cp,u}] \leftarrow$ TokenConcatenation(Solution, Compression Ratio List)
@@
To analyze the computational complexity of each component within the process described in Algorithm 3, let us consider a prompt composed of $m$ sentences, with each sentence containing $n$ tokens. The computational complexity of Algorithm 3 is broken down as follows: Line 1, Line 6 requires $\mathcal{O}(m)$; Line 3 requires $\mathcal{O}(Kmn^{2})$, where $K$ is related to the size of language model; Line 4 requires $\mathcal{O}(mn^{3})$; Line5, Line 7, and Line 9 requires $\mathcal{O}(mn)$; Line 8 requires $\mathcal{O}(m^{2}n^{2})$. In comparison, the computational complexity of Selective-Context [12] is $\mathcal{O}(Km^{2}n^{2})$, while LLMLingua [13] requires $\mathcal{O}(Kmn^{2}(1+(m-1)\tau+(m-2)\tau^{2}))$, with $\tau$ representing the compression ratio, and LLMLingua2[14] requries $\mathcal{O}(Km^{2}n^{2})$. Further technical details are elaborated in the Appendix A. Note that, under the typical conditions, $m$ and $n$ range around $20\sim 50$, $\tau$ lies between $0.2\sim 0.5$, and $K\gg m,n$. Therefore, the overall computational complexity of the PartPrompt method is $\mathcal{O}(Kmn^{2})$, resulting in the following rrelationship of computational demands among these four methods as $\mathcal{O}(\text{PartPrompt})<\mathcal{O}(\text{LLMLingua})<\mathcal{O}(\text{Selective-Context})\approx\mathcal{O}(\text{LLMLingua2})$.
@@ IV-G Theoretical Analysis
LLMLingua, another representative selective prompt compression method, consists of the budget controller, iterative token-level prompt compression, and distribution alignment [13]. The budget controller divides the prompt into paragraphs, sorting them by information entropy and preserving those with higher entropy. In PartPrompt, the node value adjustment algorithm for actual nodes is defined as $E_{\text{Adjusted}}(v_{i})\leftarrow E_{\text{Aligned}}(v_{i})\cdot M^{a_{1}}$ (Line 18 of Algorithm 1). Here $M$ is derived from the root-ward propagation and convergence of the information entropy of the corresponding segemts. In light of the fact that $M>1$ is typically the case, there exists a sufficiently large $a_{1}$ such that for any $i,j,M_{i}>M_{j}$ leads to $E_{\text{Adjusted}}(v_{i})>E_{\text{Adjusted}}(v_{j})$. In this case, parse tree compression is equivalent to retaining the part with the larger value of $M$. Therefore, the budget controller of LLMLingua can be considered as a variant of parse tree compression, which involves flattening the virtual part of the global parse tree and setting parameter $a_{1}\gg 1$ in the node value adjustment module of PartPrompt.
Considering $r_{j,i}$ as the $i$-th token of the $j$-th sentence, the iterative token-level prompt compression of LLMLingua employs Equation (11) to calculate the information entropy. This approach could be regarded as an intermediate stage between the initial version employed by Selective-Context, i.e., Equation (12), and the local approximation version of PartPrompt, i.e., Equation (13).
@@ V-A Datasets
Four representative datasets are employed to evaluate the proposed method. Concretely, BBCnews dataset is collected from articles on the BBC News website [33], and arXiv dataset is collected from papers on the arXiv preprint platform [34]. Following Selective-Context [12] and LLMLingua [13], these two datasets are used for testing contextual understanding prompts with the task of summarizing articles. To ensure that the model has not seen these data during training, we re-crawl new data for BBC News and arXiv, with all the release dates of these data occurring after Jan 1, 2024. Due to the length of arXiv dataset, which exceeds the input length that Selective-Context and LLMLingua can handle [12, 13], we truncate the first 3000 tokens of the main text, referring as truncate arXiv. The original arXiv dataset is also used to verify the applicability of our method for texts exceeding the input limit of compared methods. The PeopleDaily dataset is a dataset in Chinese constructed by crawling articles from the People’s Daily newspaper[35]. CodeNet is a code dataset. We clean a subset of CodeNet[36] to create the version used in this work. These two datasets are employed to evaluate the performance on non-English and non-natural language prompt.
There are four additional important datasets, which are tested in appendix. HotpotQA [37] is a multi-hop question answering dataset, which is designed to test the performance for question answering and multi-hop reasoning prompts. We use the same setting as in [38] and select 500 questions for testing. GSM8K [39] is a classic mathematical reasoning dataset. Following [13], we adopt the chain-of-thought prompt provided by [7] to evaluate the performance for in-context learning and chain-of-thought prompts. RULER [40] and LongBecnh [41] are two important long prompt datasets.
@@ V-B Compared Methods for Prompt Compression
As our method belongs to selective prompt compression, we compare it to the state-of-the-art selective prompt compression methods. Selective-context [12] is the first to propose selective prompt compression method. It uses a smaller LLM to calculate the information entropy of tokens and removes words with lower information entropy. The question text is fully preserved for better performance. LLMLingua [13] further proposes the model distribution alignment, budget controller and token-level iterative compression algorithm based on Selecting-context. LLMLingua receives significant improvements in context-understanding prompts, and makes selective compression method viable for chain-of-thought prompts. LLMLingua2 [14], which achieves word selective compression by distilling the compression results of GPT4 to a small model, is a completely new line approach and represents the current state-of-the-art level of prompt compression. LongLLMLingua [29] is also an important compression method, but unlike the above methods, it belongs to the question-aware approach.
@@ V-D Evaluation
Following [12, 13], we take BLEU [44], Rouge (including Rouge1, Rouge2, RougeL) [45], and BERTScore (specifically BS-F1) [46] as the metric for evaluation on BBCnews, (truncate) arXiv, PeoPleDaily and CodeNet datasets. BLEU is a composite metric of four metrics: 1-gram, 2-gram, 3-gram, and 4-gram, and is often used in the natural language processing. Rouge is classic linguistic metrics based on rule matching, while BERTScore calculates the similarity of the embeddings from BERT. The ground-truth answer is set to the output of the uncompressed prompt for the above four datasets. The evaluation for another four datasets can be found in appendix.
@@ V-E Other Settings
Apart from above settings, other experimental settings are clarified as follows. The information entropy is calculated using Llama2-7B [47], and the local parse tree is analyzed using the Stanford CoreNLP [32]. When adjusting the value in parse trees, the related hyper-parameters are set as follows: for BBCnews, $a_{1}=4$, $a_{2}=100$; for truncate arXiv, $a_{1}=3$, $a_{2}=100$; for PeopleDaily, $a_{1}=3$, $a_{2}=25$; for CodeNet, $a_{1}=1$, $a_{2}=5$. And the typical range of them are $0\leq a_{1}\leq 5$, $1\leq a_{2}\leq 1000$. The maximum token length for target LLMs is set to 300 during inference for all experiments. For the pretrained language models used in this work, Mixtral-8x7B and Llama3-70B are accessed through HuggingFace Pro API [^3] Mixtral-8x7B and Llama3-70B: https://huggingface.co/blog/inference-pro; Qwen2-72B is accessed via Aliyun or can be downloaded from HuggingFace [^4] Qwen2-72B: https://huggingface.co/Qwen/Qwen2-72B-Instruct; Llama2-7B is downloaded from HuggingFace [^5] Llama2-7B: https://huggingface.co/meta-llama/Llama-2-7b and run on an NVIDIA GeForce RTX 3090 GPU with 24G memory. More settings, e.g, using DeepSeek-LLM-7B[48] to calculate information entropy, can be found in appendix.
@@
LongLLMLingua 9.22 33.49 14.60 31.42 60.66 161.0 5.34 11.22 39.49 18.47 36.77 67.13 599.6 5.00
LLMLingua2 5.65 36.22 11.95 32.82 63.52 172.6 4.99 5.75 35.96 12.39 32.63 65.76 615.1 4.87
PartPrompt 13.69 43.54 22.87 41.05 67.58 173.7 4.96 11.73 39.83 18.74 37.01 67.32 595.8 5.03
30% tokens constraint
@@
methods PeopleDaily CodeNet
LLM Generation 10.11 36.57 15.75 33.11 67.77 795.5 2.97 7.78 36.45 14.81 33.69 63.27 820.7 1.85
20% tokens constraint
Selective-Context 3.06 26.43 8.10 23.83 61.06 477.2 4.94 4.32 29.68 10.14 27.60 58.50 341.4 4.45
LLMLingua 4.20 28.21 9.67 25.55 62.28 471.9 5.00 5.17 31.16 11.51 29.06 58.99 303.1 5.01
LongLLMLingua 6.88 32.03 12.84 29.11 64.32 473.1 4.99 4.39 30.11 10.42 27.99 58.33 303.8 5.00
LLMLingua2 3.21 27.57 8.73 24.74 61.83 504.7 4.67 2.77 27.48 8.39 25.42 57.32 310.3 4.90
PartPrompt 12.69 38.87 20.11 36.25 67.98 471.9 5.00 6.48 34.00 13.11 31.59 60.86 312.5 4.86
30% tokens constraint
Selective-Context 3.74 28.49 9.08 25.25 62.45 697.0 3.38 5.60 31.62 11.78 29.44 59.72 461.4 3.29
LLMLingua 5.81 30.92 11.56 28.09 64.33 707.3 3.33 6.62 33.05 13.24 30.84 60.02 456.9 3.33
LongLLMLingua 7.93 33.37 13.85 30.49 65.41 703.8 3.35 5.43 31.01 11.42 28.82 58.65 456.0 3.33
LLMLingua2 4.06 29.50 10.08 26.57 63.25 708.0 3.33 4.23 29.99 10.46 27.80 59.06 471.4 3.22
PartPrompt 16.39 42.29 23.34 39.48 69.97 710.0 3.32 8.29 36.70 15.50 34.30 62.33 470.2 3.23
50% tokens constraint
Selective-Context 5.07 31.02 10.86 27.85 64.09 1191.1 1.98 7.30 34.37 13.95 31.91 61.18 758.4 2.00
LLMLingua 10.34 36.84 16.44 33.71 67.85 1180.0 2.00 8.39 35.60 15.38 33.21 61.71 758.5 2.00
LongLLMLingua 11.85 37.75 17.90 34.61 67.89 1181.5 2.00 7.03 34.06 13.93 31.80 60.90 757.2 2.01
LLMLingua2 5.80 31.95 11.75 28.65 64.76 1151.3 2.05 5.65 32.35 12.23 30.07 60.03 770.3 1.97
PartPrompt 20.68 46.66 27.56 43.82 72.47 1179.1 2.00 12.05 41.95 20.04 39.40 65.59 778.0 1.95
@@ VI Experiments
Extensive experiments are conducted to demonstrate the effectiveness and superiority of PartPrompt (the proposed method) regarding various prompt compression methods, datasets, metrics, compression ratios, and scenarios. Section VI-A presents a thorough comparison between PartPrompt and other compression methods to confirm the superiority of the proposed method. Considering the effect of compression ratios and target LLMs for inference, PartPrompt is further compared with other methods for various compression ratios in Section VI-B and different target LLMs in Section VI-C. Section VI-D provides a comprehensive ablation study to verify the positive role of each component in PartPrompt. Section VI-E investigates the potential of PartPrompt under the extreme long prompt scenario. In Section VI-F and VI-E, the compressed prompt is directly compared with uncompressed one; accordingly more abundant metrics and an intuitive example are employed to show the advantages of PartPrompt in terms of text similarity and coherence maintenance.
@@ VI-A Comparative Study: Main Experiments
Comprehensive experiments are conducted in this section to fully compare the proposed method with other selective compression methods. Four diverse datasets and multiple metrics are employed to evaluate the performance. For clarity, we choose three compression ratios for analysis: 20%, 30%, and 50%. As the actual compression ratio is not under the control of some methods, the average length of the compressed prompt (named as “tokens”) and inverse compression ratio ($1/\tau$) are also listed to show the actual compression result. The LLM Generation is presented with a single compression ratio, since there is no compression ratio to set.
The experimental results are shown in Table I. Overall, it is evident that PartPrompt achieves significant improvements across all datasets, compression ratios, and metrics, reaching state-of-the-art (SOTA) performance. Specifically, PartPrompt substantially outperforms the baseline methods in terms of BLEU and ROUGE scores. Notable gains are also observed on BERT scores with PartPrompt. In addition, the LLMLingua2 method, being the second generation of the LLMLingua series, has not made an improvement compared to LLMLingua. This is because LLMLingua2 is not, in the conventional sense, an upgraded version of LLMLingua; rather, it represents an entirely novel approach and is good at other scenarios.
Apart from comparing to selective compression methods, PartPrompt is further compared to the generative compression method. Qwen2-72B is employed for generating compressed prompts, which is denoted as LLM Generation in Table I. Note that, using an LLM for compression needs a significant amount of computational resources, which is contrary to the original goal of prompt compression and therefore impractical. Despite that, LLM Generation may still serve as a reference method. These results show that PartPrompt outperforms LLM Generation on all four datasets, which not only proves the advantage of PartPrompt but also validates the rationality of selective compression methods.